# Evaluating the Instruction-following Abilities of Language Models using Knowledge Tasks

### v1

@@ Abstract
In this work, we focus our attention on developing a benchmark for instruction-following where it is easy to verify both task performance as well as instruction-following capabilities. We adapt existing knowledge benchmarks and augment them with instructions that are a) conditional on correctly answering the knowledge task or b) use the space of candidate options in multiple-choice knowledge-answering tasks. This allows us to study model characteristics, such as their change in performance on the knowledge tasks in the presence of answer-modifying instructions and distractor instructions. In contrast to existing benchmarks for instruction following, we not only measure instruction-following capabilities but also use LLM-free methods to study task performance. We study a series of openly available large language models of varying parameter sizes (1B-405B) and closed source models namely GPT-4o-mini, GPT-4o. We find that even large-scale instruction-tuned LLMs fail to follow simple instructions in zero-shot settings. We release our dataset, the benchmark, code, and results for future work.
@@ 1 Introduction
The growth of increasingly powerful large language models has resulted in the development of end-user applications including assistants for coding and software engineering (Ozkaya, 2023; Zhang et al., 2023; Ross et al., 2023), workflow and business automations (Grohs et al., 2023; Wornow et al., 2024), self-help assistants (Zhou et al., 2020; Shuster et al., 2022) and more. The need for highly accurate and controllable systems that follow precise instructions have led to the development of methods to improve reliability and consistency in the output for LLMs. Such methods include few-shot prompting (Gao et al., 2020; Kojima et al., 2022b), reasoning with explanations (Wei et al., 2022; Huang & Chang, 2022), checking for consistency/self-consistency (Wang et al., 2022), use of intermediate evaluators or LLMs operating as judges (Zheng et al., 2023), etc.
While there has been a lot of focus on assessing the knowledge of LLMs (Brown et al., 2020; Heinzerling & Inui, 2021), logical reasoning (Hendrycks et al., 2021; Wei et al., 2022; Ma et al., 2024), programmatic ability (Dakhel et al., 2023; Chen et al., 2021), problem solving ability (Lightman et al., 2024), etc, the study of their ability to follow precise instructions is relatively nascent; works such as FoFo (Xia et al., 2024), InFoBench (Qin et al., 2024), RuleBench  (Sun et al., 2024), IFEval (Zhou et al., 2023b) attempt to address this gap. While FoFo  Xia et al. (2024) assesses the ability of models to generate outputs conforming to existing real-world output formats such as the HL7-CDA format used in Healthcare applications, RuleBench (Sun et al., 2024) assesses a model’s capabilities on inferential rule-following using rules which can be encoded in instructions and first-order-logic. On the other hand, benchmarks such as InFoBench (Qin et al., 2024) and IFEval (Zhou et al., 2023b) assess the ability of LLMs to follow arbitrary task specific instructions though neither InFoBench nor IFEval provide easy ways of verifying (i) task success and (ii) instruction following capabilities simultaneously (see Section 2 for a detailed discussion).
In this work, we focus our attention on developing a benchmark for instruction-following where it is easy to verify both task performance as well as instruction following capabilities. We adapt existing commonly used knowledge benchmarks including MMLUPro (Wang et al., 2024a), MathQA (Amini et al., 2019), Winogrande (Sakaguchi et al., 2021), BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020) and augment them with two broad classes of instructions: (i) Instructions that are conditional on the answer to the question (ii) Instructions that are applied uniformly regardless of the answer or task. We include a detailed study of multiple LLMs and find that even the largest models have trouble following relatively simple instructions.
Our contributions are as follows: (i) We release the first benchmark that assesses the zero-shot instruction-following performance of models using knowledge and reasoning question-answering (QA) tasks. (ii) We employ multiple QA-conditioned instructions to examine instruction-following performance across different instruction classes, including those dependent on answer-type. (iii) We include instruction instances that serve as distractors for the original knowledge-tasks (iv) Unlike previous studies, we use LLM-free evaluation metrics to assess both knowledge and instruction-following abilities. (v) We offer automated error analysis measures, pre-classifying likely errors for each instruction instance. (vi) Our benchmark creation method is easy to extend to new instructions and datasets.
@@ 2 Related Work
Table 1: Comparison with existing Instruction Following benchmarks
Benchmark Task
Deterministic
Outputs
Content
Verification
QA-Conditioned
Instructions
Evaluator
FoFo Eval (Xia et al., 2024) Format following LLM
IFEval (Zhou et al., 2023b) Instruction Following Direct
InFoBench (Qin et al., 2024) Instruction Following LLM
RuleBench (Qin et al., 2024) Inferential Rule Following Direct
This work Instruction Following Direct
@@
Evaluating the capabilities of large language models (LLMs) has been a significant area of research, with studies focusing on various aspects of LLM performance. Researchers have developed multiple benchmarks to assess factual knowledge (Petroni et al., 2019; Roberts et al., 2020; Lin et al., 2022), logical reasoning abilities (Wei et al., 2022; Zhou et al., 2023a; Saparov et al., 2023), general problem-solving capabilities (Kojima et al., 2022a) and more.
@@
Recently there have also been studies on instruction-following - for instance, FoFo (Xia et al., 2024) evaluates models on format-following tasks and studies the ability of LLMs to generate outputs in existing real-world formats. In a similar vein, IFEval (Zhou et al., 2023b) assesses LLMs’ ability to follow arbitrary task-specific instructions (e.g.) based on response length, casing, etc, focusing primarily on whether the instructions are followed rather than the correctness of the output for the task. InFoBench (Qin et al., 2024) advances this research by introducing a metric known as the ‘Decomposed Requirements Following Ratio’ (DRFR) which is based on each aspect of an instruction that needs to be met. Along with $500$ diverse instructions and $2,250$ decomposed questions, InFoBench offers performance evaluation using OpenAI’s GPT4, across multiple constraint categories and highlights key areas where advanced LLMs can improve in complex instruction-following tasks. LLMBar (Zeng et al., 2024) contributes to this area by providing a meta-evaluation benchmark specifically designed to test an LLM evaluator’s ability to discern instruction-following outputs. The benchmark consists of $419$ manually curated pairs of outputs, where one output adheres to instructions and the other, while potentially more engaging or deceptive, does not.
Recent research also acknowledges the growing body of work focused on the ability of LLMs to handle multiple instructions concurrently, as explored in the Multi-Task Inference Benchmark (Son et al., 2024). This benchmark reveals that LLMs can perform better when performing multi-task inference than single-task inference. In addition to this, the FollowEval(Jing et al., 2023) benchmark addresses the importance of instruction-following capacity by providing a dual-language assessment crafted by human experts, covering various reasoning dimensions. Moreover, the FollowBench(Jiang et al., 2024) benchmark emphasizes fine-grained constraint evaluation, utilizing a multi-level mechanism to assess LLM performance across different types of constraints. Lastly, the CELLO(He et al., 2024) benchmark evaluates LLMs’ capability to comprehend complex instructions, identifying limitations in existing assessments and offering a systematic framework for evaluation.
@@
Our work builds upon these efforts by developing a benchmark that allows for easy verification of both task performance and instruction-following capabilities simultaneously. We augment existing knowledge benchmarks by creating instructions that are conditional on answering the QA-based knowledge task correctly. We also include instructions that are applied on the candidate space of answers provided in these knowledge tasks. Our approach of applying instructions on knowledge tasks provides an easy way of measuring performance. Further, it also allows us to study the interactions between knowledge and instruction following, and to investigate whether instructions serve as distractors for the original knowledge task when the instructions should result in no change to the original answer of the knowledge task.
3 Instruction-Following Evaluation Dataset
We now describe the process for creating our evaluation dataset.
Design Principles: We develop our instructions keeping the following design principles in mind: (i) We would like instructions to be unambiguous and be presented in a way that can be communicated clearly - if humans cannot follow the instructions and agree on the same output, LLMs should and likely would not be able to. (ii) We would like them to be easy to follow and not require complex reasoning abilities to follow so that models at all scales have a fair chance of success, (iii) The instructions need to have deterministic outputs that use the original answers of the knowledge-task or the candidate space of answers, or both, so that they can be evaluated easily with instruction specific scorers. (iv) We would like our benchmark to be based on a diverse mix of knowledge tasks, and be easily extensible to new ones.
3.1 Knowledge and Reasoning Tasks
(i) MMLUPro (Wang et al., 2024a): MMLUPro extends the MMLU dataset to make it more challenging by a) increasing the number of options from four to ten and b) increasing problem difficulty by focusing on more reasoning oriented problems. We consider all $14$ subjects in the MMLUPro benchmark. We cap the maximum number of samples for each subject to be $150$ samples.
We also select a few common-sense and reasoning datasets:
3.2 Instruction Categories
@@
Operations on List use_options_to_create_string
Use the last character of the text associated with
every candidate answer to create a string
7867 950
sort_options_to_create_string
@@
Unlike datasets that require open-ended generation for answering, our selected tasks have a structured answer-space. This allows us to craft instructions using these answer-spaces in a way that can be verified easily. We define the following instruction categories.
(i) String Manipulation: This operation involves manipulating the characters within the correct answer. We apply simple transformations like changing the case of the answer text or reversing the answer text, etc.
(ii) Format Correct Answer: This operation involves displaying the correct answer in the specified format. This involves printing any numeric answers in words or appending a string to the correct answer, etc.
(iii) Numeric Manipulation: This instruction involves incrementing a numeric quantity by one and has no effect on non-numeric answer text.
(iv) Operations on list (Conditional): This operation involves conditionally manipulating the candidate answer space – for instance, incrementing incorrect answers by one, sorting the incorrect answers, etc.
(v) Operations on list: These are simple instructions that do not depend on the correct answer of the original knowledge tasks. Examples, include - sorting all candidate options, concatenating characters from each candidate option, etc.
For each instruction category, we create multiple instructions. Table 2 presents the $13$ instructions we have included in our work. The task prompts (instructions) for each of the $13$ instruction types with an example are available in the Appendix (Section A.1).
Instruction Creation: To create each instruction, the authors iteratively refined them until all the authors had complete agreement in the output when they followed them manually. Examples of aspects of iterative improvement include - explicitly making clear what is not to be included in the output, how the output is to be presented, etc. We then asked $2$ computer science researchers to follow and generate the output for $75$ instructions across all our instruction types and datasets. We found that both the researchers were able to follow our instructions successfully and generated the same response for $93.33\%$ of the instances. The first annotator generated the correct response for $98.67\%$ of the instances, while the second annotator for $94.67\%$ of the instances. Upon analyzing their responses, we found the only instruction-following error was rounding off the decimal number when truncating to two decimal places. We also found very few human errors in the annotator’s response, specifically for instructions like $reverse\_correct\_answer\_alternate\_case$ on datasets with long output text such as PIQA.
@@
Answering baseline-instructions: We additionally develop two baseline instructions – (1) printing the correct answer option [^1] We use ‘label’ and ‘option’ interchangeably to denote the candidates in a multiple-choice QA task. from the candidate space ($print\_correct\_answer\_label$), and (2) printing only the text associated with the correct answer option ($print\_correct\_answer\_text$).
@@
Instructions with no-effect: Certain instructions may be inapplicable for some knowledge tasks. For example, in the MathQA dataset, some instances have none of these as the correct answer and are not numeric. Here, instructions such as $numformat\_numeric\_answer$ or $increment\_correct\_numeric\_answer\_by\_one$ will not affect the existing answer of the knowledge-task. We refer to these instructions as “distractor” instances and expect that in these instances, models should perform as well as they do on the answering baseline-instructions. We include details and statistics of such instructions in the Appendix (Section A.3 Tables 7 and 9).
@@ 3.3 Automated classification of errors
KnowledgeError: For every instruction instance, this set includes: (i) Answering with the incorrect candidate answer – regardless of whether it is its label or text, (ii) Applying the instruction (when applicable) on the incorrect candidate answer.
Note that these error sets are not mutually exclusive – a model response to the alternate casing instruction that starts with lowercase and is applied on the incorrect answer candidate is an ‘IFError’ as well as a ‘KnowledgeError’. Further, these error sets are meant to be high-precision to aid analysis and cannot reasonably expected to anticipate and cover everything LLMs might respond with. Hence, there will be model errors that we cannot classify and we intend to continue to expand the benchmark’s error classification capabilities. We would also like to highlight that if models make a lot of errors that are not classified, it is likely that those are more severe instruction-following errors.
3.4 Metrics
Exact Match: We report the model performance as exact match under two settings - strict and loose. In the strict setting, we perform basic string parsing (removing beginning and ending whitespaces, quotations, etc.) and compare the model prediction to the expected output for the applied instruction.
However, we observe that models often make errors when following the primary instruction. These could be minor copying errors, such as missing a period or comma, or even fixing typos within the provided options. On the other hand, they could also be instruction following mistakes, where for instance the option label is added to the response even when the prompt explicitly states otherwise. Given that we do not expect models to make such mistakes given clear instructions, we use the strict metric in the majority of our evaluations.
However, we also define a relaxed version of the exact match called loose exact match, allowing for a Levenshtein distance (Levenshtein, 1966) of two edit operations between the prediction and ground truth. Additionally, we also perform whitespace-free matching as part of our loose criterion. Similar to Zhou et al. (2023b), we consider our loose match as a complement to the strict one.
@@ 3.5 Benchmark Dataset
Full Benchmark: We select a subset of $1500$ samples randomly from each datasets and apply each applicable instruction on the same. For MMLUPro, we consider a subset of $150$ samples per subject and apply each applicable instruction.
Lite Benchmark: We select a subset of $150$ samples randomly from the full version created above for each dataset and apply each applicable instruction on the same. For MMLUPro, we consider a subset of $25$ samples per subject and apply each applicable instruction. Statistics for the above two versions are available in presented in Table 2. Detailed statistics for each dataset and the instruction types are provided in the appendix section A.3. Additionally, each benchmark includes a set of instances when instructions have no effect (called the no-effect subset).
3.5.1 Benchmark ranking
4 Evaluation
We present an evaluation on our benchmark using a variety of models and study the following research questions: (i) Do models display a difference in performance on the two simple baseline instruction tasks? (ii) Do models display a variation in performance across our different instruction categories? (iii) Are models robust to, or get distracted by instructions that do not apply to the task? (iv) Does the size of a model impact its instruction-following capability?
4.1 Models and Inference
Table 3: List of Models evaluated on our benchmark. For ease of reference, we use short names by considering only the model name and the parameter size
@@ 4.1 Models and Inference
We evaluate our benchmark on a range of open instruction-tuned models and parameter sizes. For ease of presentation, we categorize them based on their parameter count as shown in Table 3. Our inference code uses vLLM (Kwon et al., 2023) for running the evaluations. We use greedy decoding for generations and bf16 as floating point precision. We generate a maximum of 1024 tokens per instance. We use A100 80GB GPUs for running inference. We use an instance hosted by a cloud provider for Llama-3.1-405B-Instruct, while we use OpenAI APIs for GPT4-o and GPT4-o-mini models.
In all our experiments, we perform zero-shot Chain-of-Thought (CoT) (Wei et al., 2024) reasoning. Models see the same prompt based on prompt guides for the original knowledge tasks in lm-evaluation-harness framework (Gao et al., 2024) and OpenAI evals. [^2] https://github.com/openai/simple-evals We instruct the model to generate reasoning first and then the answer (See examples in Appendix Section A.1). We write custom post-processing scripts to extract the model’s answer as described in the next section.
4.2 Output Post-processing
All our task prompt templates, as shown in Appendix A.1, explicitly instruct the model to provide their final response after a ‘Response:’ keyword. As part of our strict evaluation metric (Section 3.4), we search for and extract the response after this keyword while computing the metrics.
[…]
Figure 1: Left: Average exact match performance across all tasks for the print_correct_answer and print_correct_answer_label instructions. Right: Knowledge and instruction following (IF) errors across all tasks for the print_correct_answer instruction. A lower error is better. Both results shown using Full Benchmark data. Lite Benchmark results can be found in Appendix Figure 6.

### v2

@@ Abstract
LLM evaluation benchmarks have traditionally separated the testing of knowledge/reasoning capabilities from instruction following. In this work, we study the interaction between knowledge and instruction following, and observe that LLMs struggle to follow simple answer modifying instructions, and are also distracted by instructions that should have no bearing on the original knowledge task answer. We leverage existing multiple-choice answer based knowledge benchmarks and apply a set of simple instructions which include manipulating text (eg.: change case), numeric quantities (eg.: increase value, change formatting), operate on lists (eg.: sort answer candidates) and distractor instructions (eg.: change case of numeric answers).
We evaluate models at varying parameter sizes (1B-405B) from different model families and find that, surprisingly, all models report a significant drop in performance on such simple task compositions. While large-sized and frontier models report performance drops of 40-50%, in small and medium sized models the drop is severe (sometimes exceeding 80%). Our results highlight a weakness in the traditional separation of knowledge/reasoning and instruction following, and suggest that joint-study of these capabilities are important. We release our benchmark dataset, evaluation framework code, and results for future work.
Evaluating the Instruction-following Abilities of Language Models using Knowledge Tasks
Rudra Murthy IBM Research AI rmurthyv@in.ibm.com                        Praveen Venkateswaran IBM Research AI praveen.venkateswaran@ibm.com
Prince Kumar IBM Research AI prince.kumar12@ibm.com                        Danish Contractor IBM Research AI danish.contractor@ibm.com
@@ 1 Introduction
The growth of increasingly powerful large language models (LLMs) has resulted in the development of end-user applications including assistants for coding and software engineering (Ozkaya, 2023; Zhang et al., 2023; Ross et al., 2023), workflow and business automations (Grohs et al., 2023; Wornow et al., 2024), self-help assistants (Zhou et al., 2020; Shuster et al., 2022) and more. The need for highly accurate and controllable systems that follow precise instructions have led to the development of datasets and methods to improve reliability and consistency in the output for LLMs. Such methods include few-shot prompting (Gao et al., 2020; Kojima et al., 2022a), reasoning with explanations (Wei et al., 2022; Huang and Chang, 2022), checking for consistency/self-consistency (Wang et al., 2022), use of intermediate evaluators or LLMs operating as judges (Zheng et al., 2023), etc.
Figure 1: Average exact match performance across all tasks for the print_correct_answer (PCA) and print_correct_answer_label (PCA Label) instructions.
However, despite such rapid progress and benchmark saturation we find that LLMs struggle on very simple tasks. For instance, as shown in Figure 1, given a multiple-choice question with option labels and their text, if we instruct the models to print the text associated with the correct answer instead of the answer label, we observe a significant drop ($\sim 20\%$ on average) in knowledge-task performance.
From a knowledge perspective, this task is no harder than selecting (generating) the answer label. Yet, this pattern is consistent even for frontier models like GPT-4o (Appendix Figure 5). We hypothesize that this drop in performance could, in part, be due to the training process resulting in models being over-fit to certain input/output task formats, resulting in worse instruction following for other formats. However this cannot explain other observations – for instance, including additional distractor instructions that have no effect on the answering task, results in a further 5-30% drop (Section 4.2.3).
In this paper, we study this interaction of knowledge and instruction following; real-world use of LLMs typically involves both aspects, and evaluation benchmarks which specifically focus on one aspect or the other may not reflect true model performance. To keep evaluation easy and verifiable, our tasks are based on multiple-choice answers instead of long-form text generation tasks, which are harder to evaluate reliably for joint knowledge and instruction-following capabilities.
We adapt commonly used knowledge and reasoning benchmarks including MMLUPro (Wang et al., 2024), MathQA Amini et al. (2019), Winogrande (Sakaguchi et al., 2021), BoolQ (Clark et al., 2019), PIQA Bisk et al. (2020) and augment them with two broad classes of instructions: (i) Instructions that are conditional on the answer to the question (ii) Instructions that are applied uniformly regardless of the answer or task. Our instructions are simple and include those that manipulate the text (eg.: change case), numeric quantities (eg: increase value, change formatting), operate on lists (eg: sort answer candidates). Our choice of using multiple-choice answers-based tasks with simple answer-modifying instructions, also allows us to automate error analysis for commonly occurring error-types and classify errors based on their origin - knowledge/reasoning errors vs instruction-following errors. Our results show that we can automatically classify approximately $80\%$ of errors.
Contributions: In summary, we make the following contributions: (i) We present an evaluation framework that supports LLM-free scoring to study knowledge-conditioned instruction-following, (ii) Our experimental design enables automated error analysis (to the extent that LLMs dont have catastrophic generation failures), (iii) We create an evaluation benchmark using $5$ popular datasets, consisting of $13$ instruction types across $5$ different categories. It can be easily extended to additional tasks and instructions, (iv) We use this benchmark to evaluate $19$ models at varying parameter sizes (1B-405B) from different model families and find that, surprisingly, all models studied report a significant drop in performance on simple single-instruction task composition. While large-sized and frontier models report performance drops of 40-50%, in small and medium sized models the drop is severe (sometimes exceeding 80%). We also observe that, even within a model family, a larger parameter model may perform comparably to its smaller sibling on standard knowledge/reasoning tasks, but smaller ones drop more significantly on in our evaluation tasks. We release all code and data to enable future work.
@@ 2 Related Work
Evaluating the capabilities of large language models (LLMs) has been a significant area of research, with studies focusing on various aspects of LLM performance. Researchers have developed multiple benchmarks to assess factual knowledge Petroni et al. (2019); Roberts et al. (2020); Lin et al. (2022), logical reasoning abilities Wei et al. (2022); Zhou et al. (2023a); Saparov et al. (2023), general problem-solving capabilities Kojima et al. (2022b) and more.
@@
Recently there have also been studies on instruction-following - for instance, FoFo Xia et al. (2024) evaluates models on format-following tasks and studies the ability of LLMs to generate outputs in existing real-world formats. In a similar vein, IFEval Zhou et al. (2023b) assesses LLMs’ ability to follow arbitrary task-specific instructions (e.g.) based on response length, casing, etc, focusing primarily on whether the instructions are followed rather than the correctness of the output for the task. InFoBench Qin et al. (2024) advances this research by introducing a metric known as the ‘Decomposed Requirements Following Ratio’ (DRFR) which is based on each aspect of an instruction that needs to be met. Along with $500$ diverse instructions and $2,250$ decomposed questions, InFoBench offers performance evaluation using OpenAI’s GPT4, across multiple constraint categories and highlights key areas where advanced LLMs can improve in complex instruction-following tasks. LLMBar (Zeng et al., 2024) is another contribution to this area, as it provides a meta-evaluation benchmark specifically designed to test an LLM evaluator’s ability to discern instruction-following outputs. The benchmark consists of $419$ manually curated pairs of outputs, where one output adheres to instructions and the other, while potentially more engaging or deceptive, does not. ComplexBench (Wen et al., 2024) assesses the ability of LLMs to follow arbitrary task specific instructions, while Meowpass (He et al., 2024) presents a benchmark to evaluate complex instruction following by dynamically composing multiple IFEval style instructions, and consists of $1500$ diverse prompt and instruction pairs. Another study introduces ManyIFEval (Harada et al., 2025), a benchmark assessing LLMs’ ability to follow multiple instructions, revealing a “curse of instructions” where accuracy declines as instruction count increases. Li et al. (2024) propose a method to evaluate instruction following ability via verbalizer manipulation. Specifically, they modify the classification task labels with different verbalizers which may or may not be semantically relevant to the task. They observe that all models fail to follow instructions when they instruct the model to flip the labels (unnatural setting).
@@
Our work complements these efforts by developing a benchmark that allows for easy verification of both task performance and instruction-following capabilities simultaneously. We augment existing knowledge benchmarks by creating instructions that are conditional on answering the QA-based knowledge task correctly. We also include instructions that are applied on the candidate space of answers provided in these knowledge tasks. Our approach of applying instructions on knowledge tasks provides an easy way of measuring performance as well as automating error analysis. Further, it also allows us to study the interactions between knowledge and instruction following, and to investigate whether instructions serve as distractors for the original knowledge task when the instructions should result in no change to the original answer of the knowledge task.
3 Dataset Creation
@@
Operations on List (OOL) sort_options_to_create_string
@@
We now describe our process for creating our evaluation dataset.
3.1 Knowledge and Reasoning Tasks
Our framework supports including any knowledge task with a fixed answer space - we select the following commonly used benchmarks: MMLUPro  (Wang et al., 2024), MathQA (Amini et al., 2019), BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), and Winogrande (Sakaguchi et al., 2021) as the basis for our knowledge-grounded instruction-following benchmark. These datasets involve either binary classification or multiple-choice-questions (MCQs) spanning different reasoning and problem-solving skills. We select $1500$ samples randomly from each dataset and apply answer-modifying instructions as described in the next section.
3.2 Instruction Categories
@@
Unlike datasets that require open-ended generation for answering, our selected tasks have a structured answer-space. This allows us to craft instructions using these answer-spaces in a way that can be verified easily. We summarize our instruction categories in Table 1. The task prompts (instructions) for each of the $13$ instruction types with an example are available in the Appendix (Section B.4).
Instruction Creation: To create each instruction, the authors iteratively refined them until all the authors had complete agreement in the output when they followed them manually. Examples of aspects of iterative improvement include - explicitly making clear what is not to be included in the output, how the output is to be presented, etc. We include additional details of the instruction-writing process in the Appendix (B.2).
@@
Answering baseline-instructions: Given the multiple-choice answer knowledge benchmarks, we consider two baseline instructions – (1) printing the correct answer option [^1] We use ‘label’ and ‘option’ interchangeably to denote the candidates in a multiple-choice QA task. from the candidate space ($print\_correct\_answer\_label$), and (2) printing only the text associated with the correct answer option ($print\_correct\_answer\_text$).
@@
Instructions with no-effect: Certain instructions may be inapplicable for some knowledge tasks. For example, in the MathQA dataset, some instances have none of these as the correct answer and are not numeric. Here, instructions such as $numformat\_numeric\_answer$ or $increment\_correct\_numeric\_answer\_by\_one$ will not affect the existing answer of the knowledge-task. We refer to these instructions as “distractor” instances and expect that in these instances, models should perform as well as they do on the original answering task. We include details and statistics of such instructions in the Appendix (Section B.5 Tables 7 and 9).
3.3 Metrics and Output Post-processing
We report the model performance as a form of exact match (EM) where we perform basic string parsing (removing beginning and ending whitespaces, quotations, etc.) and compare the model prediction to the expected output for the applied instruction.
[…]
However, models do still make errors when following the primary instruction. These could be minor copying errors, such as missing a period or comma, or even fixing typographical errors within the provided options. On the other hand, they could also be instruction following mistakes, where for instance, the option label is added to the response even when the prompt explicitly states otherwise. We do not expect models to make such mistakes given clear instructions, and therefore, penalize models for these types of errors in our evaluation.
@@ 3.4 Automated classification of errors
KnowledgeError: For every instruction instance, this set includes: (i) Answering with the incorrect candidate answer – regardless of whether it is its label or text, (ii) Applying the instruction correctly (when applicable) on the incorrect candidate answer.
Note that these error sets are not mutually exclusive – a model response to the alternate casing instruction that starts with lowercase and is applied on the incorrect answer candidate is an ‘IFError’ as well as a ‘KnowledgeError’. Further, these error sets are meant to be high-precision to aid analysis and cannot reasonably expected to anticipate and cover everything LLMs might respond with. Hence, there will be model errors that we cannot classify [^3] Errors are matched using the same output-processing used for ground-truth scoring. and we intend to continue to expand the benchmark’s error detection capabilities. We would also like to highlight that if models make a lot of errors that are not classified, it is likely that those are more severe instruction-following errors.
@@ 3.5 Benchmark Dataset
Full Benchmark: We select a subset of $1500$ samples randomly from each dataset and apply each applicable instruction on the same. For MMLUPro, we consider a subset of $150$ samples per subject and apply each applicable instruction.
Lite Benchmark: We select a subset of $150$ samples randomly from the full version created above for each dataset and apply each applicable instruction on the same. For MMLUPro, we consider a subset of $25$ samples per subject and apply each applicable instruction. Statistics for the above two versions are available in presented in Table 1. Detailed statistics for each dataset and the instruction types are provided in the appendix section B.5. Additionally, each benchmark includes a set of instances when instructions have no effect (called the no-effect or distractor subset).
@@
Table 2: List of Models evaluated on our benchmark.
3.5.1 Benchmark ranking
4 Evaluation
We present an evaluation on our benchmark using a variety of models and study the following research questions: (i) Do models display a difference in performance on the two simple answering baseline-instruction tasks? (ii) Do models display a variation in performance across our different instruction categories? (iii) Are models robust to, or get distracted by instructions that do not apply to the task? (iv) Does the size of a model impact its instruction-following capability?
4.1 Models and Inference
We evaluate a range of open instruction-tuned models and parameter sizes using our benchmark. For ease of presentation, we categorize them based on their parameter count as shown in Table 2. Our inference code uses vLLM Kwon et al. (2023) for running the evaluations. We use greedy decoding for generations and bf16 as floating point precision. We generate a maximum of 1024 tokens per instance. We use A100 80GB GPUs for running inference. We use an instance hosted by a cloud provider for Llama-3.1-405B-Instruct, while we use OpenAI APIs for GPT4-o and GPT4-o-mini models.
In all our experiments, we perform zero-shot Chain-of-Thought (CoT) Wei et al. (2024) reasoning. Models see the same prompt based on prompt guides for the original knowledge tasks using the lm-evaluation-harness framework Gao et al. (2024) and OpenAI evals. [^4] https://github.com/openai/simple-evals We instruct the model to generate reasoning first and then the answer (See examples in Appendix Section B.4). We write custom post-processing scripts to extract the model’s answer as described previously.
Model PCA label PCA String Numeric Format OOL Conditional OOL Avg. Drop
Qwen2.5-1.5B-Instruct 0.36 -0.20 -0.35 -0.34 -0.32 -0.29 -0.35 -0.33 / -90.5%
Qwen2.5-3B-Instruct 0.44 -0.12 -0.39 -0.42 -0.2 -0.3 -0.24 -0.31 / -69.7%
Qwen2.5-7B-Instruct 0.55 0.11 -0.41 -0.11 -0.33 -0.32 -0.38 -0.31 / -56.5%
Qwen2.5-14B-Instruct 0.75 -0.20 -0.60 -0.21 -0.57 -0.30 -0.29 -0.39 / -52.8%
Qwen2.5-32B-Instruct 0.79 -0.01 -0.47 -0.17 -0.40 -0.27 -0.26 -0.31 / -39.9%
Qwen2.5-72B-Instruct 0.81 -0.04 -0.50 -0.25 -0.31 -0.27 -0.18 -0.30 / -37.3%
Llama-3.2-1B-Instruct 0.10 -0.01 -0.08 -0.08 -0.09 -0.07 -0.10 -0.08 / -87.6%
Llama-3.2-3B-Instruct 0.47 -0.18 -0.37 -0.37 -0.36 -0.37 -0.45 -0.38 / -81.5%
Llama-3.1-8B-Instruct 0.59 -0.09 -0.47 -0.31 -0.37 -0.42 -0.45 -0.40 / -67.7%
Llama-3.1-70B-Instruct 0.79 -0.06 -0.42 -0.39 -0.41 -0.41 -0.43 -0.41 / -52.5%
Phi-3.5-mini-instruct 0.48 -0.14 -0.43 -0.33 -0.35 -0.37 -0.34 -0.36 / -75.5%
Phi-3-small-8k-instruct 0.58 -0.05 -0.47 -0.43 -0.34 -0.42 -0.52 -0.44 / -74.8%
Phi-3-medium-4k-instruct 0.72 -0.07 -0.53 -0.50 -0.46 -0.37 -0.40 -0.45 / -63.2%
Gemma-2-9b-it 0.62 -0.05 -0.50 -0.37 -0.29 -0.30 -0.30 -0.35 / -56.5%
Gemma-2-27b-it 0.69 -0.08 -0.44 -0.37 -0.30 -0.23 -0.22 -0.31 / -45.3%
GPT4-o-mini 0.74 -0.14 -0.39 -0.36 -0.35 -0.31 -0.30 -0.34 / -45.8%
GPT4-o 0.81 -0.04 -0.32 -0.32 -0.37 -0.28 -0.18 -0.29 / -36.2%
Llama-3.1-405B-Instruct 0.83 -0.07 -0.39 -0.41 -0.41 -0.39 -0.45 -0.34 / -41.0%
Table 3: As compared to printing the correct answer label (PCA Label) which serves as the baseline for the vanilla knowledge task, printing the correct answer (PCA) results in a significant drop in exact match performance. Applying additional instructions on the answer space of the questions result in an even more severe drop (on every category). Drop in performance reported as compared to PCA Label. Frontier model results reported on Lite Benchmark data.
Figure 2: Knowledge and instruction following (IF) errors across all tasks for the print_correct_answer instruction. A lower error is better. Results shown using Full Benchmark data. Lite Benchmark results can be found in Appendix Figure 5.